Twenty foxes are captured in a forest, tagged, then released. Several weeks later, 20 foxes are captured in the same forend 8 are found to have tags.Use this information to estimate how many foxes are in the forest.
Answer by stanbon(48558)

Twenty foxes are captured in a forest, tagged, then released. Several weeks later, 20 foxes are captured in the same forend 8 are found to have tags.Use this information to estimate how many foxes are in the forest.
---------
The 20 tagged foxes become a new, evenly distributed, population in the forest.
You found that when you captured 20 from the general population, 8 were tagged.
EQUATION:
Let the size of the original population be "x"
20/x = 8/20
Invert both sides to get:
x/20 = 20/8
x=20(20/8)
x=400/8
x=50 (number of foxes in the forest)
Cheers,
Stan H.
.